# Learning to Plan for Retrieval-Augmented Large Language Models from Knowledge Graphs

### v1

@@ Abstract
Junjie Wang11footnotemark: 1, Mingyang Chenthanks:   Equal contribution., Binbin Hu, Dan Yang, Ziqi Liu, Yue Shen, Peng Wei, Zhiqiang Zhang, Jinjie Gu, Jun Zhou, Jeff Z. Pan, Wen Zhang22footnotemark: 2, Huajun Chenthanks:   Corresponding authors. Zhejiang University, Ant Group, The University of Edinburgh Zhejiang University - Ant Group Joint Laboratory of Knowledge Graph {wangjj2018,mingyangchen,zhang.wen,huajunsir}@zju.edu.cn https://github.com/zjukg/LPKG
@@ 1 Introduction
The past few years have witnessed significant innovations in LLMs Ouyang et al. (2022); Touvron et al. (2023); Chowdhery et al. (2023); AI@Meta (2024). While LLMs excel in many natural language processing tasks, they still face challenges, particularly the smaller models, in handling complex question-answering (QA) tasks Press et al. (2023); Shao et al. (2023); Yao et al. (2022).
@@
To improve the performance of LLMs on complex QA tasks, past research has tried various methods: (1) Employing carefully designed prompt strategies to guide the model in reasoning, such as Chain of Thought (CoT) Kojima et al. (2022); Wei et al. (2022) and Tree of Thought (ToT) Yao et al. (2024) methods; (2) Utilizing retrieval techniques to obtain supplemental information from external knowledge source Lewis et al. (2020); Guu et al. (2020); (3) Combining prompt strategies with retrieval enhancements, as exemplified by methods like ReAct Yao et al. (2022) and Self-Ask Press et al. (2023). The third approach has garnered widespread research interest due to its integration of the advantages of the first two methods. The fundamental idea of this class of methods is to guide LLMs in breaking down a complex question into multiple simpler sub-questions and then use a retrieval-augmented generation (RAG) method to answer each sub-question, thereby deducing the answer to the original complex question. However, planning for complex questions is non-trivial, especially for smaller LLMs (with fewer than 10 billion parameters), which often require supervised fine-tuning Aksitov et al. (2023); Chen et al. (2023a); Qin et al. (2023).
@@
KGs usually store accurate knowledge in a structured way. We find that a KG pattern can be viewed as the abstract of a complex question, as shown in Figure 1, which reveals the connection between question planning and patterns. This opens up the possibility of constructing training data to enhance the planning capabilities of LLMs using KGs. Specifically, we start by grounding predefined patterns in an open-domain KG to extract numerous instances, which we then verbalize into complex questions and corresponding sub-questions in natural language. In this way, we effectively create a large number of accurate planning data for fine-tuning. Being fine-tuned with these planning data, LLMs’ capability of generating plans for complex questions is enhanced, resulting in better final answers by parsing and executing these plans. We refer to this innovative framework as Learning to Plan from Knowledge Graphs (LPKG).
@@ 2 Related Works
In the existing realm of LLMs, KGs are primarily utilized as sources of structured factual knowledge. For example, Think-on-Graph Sun et al. (2023) extracts relevant triples from KGs to assist in QA. Reasoning on Graph (RoG) Luo et al. (2023) generates relation-based plans and retrieves corresponding paths from these graphs. While aiding in KGQA tasks where answers are directly sourced from KGs, these graphs also support rationale generation. Chain-of-Knowledge (CoK) Li et al. (2024) further leverages KGs along with other heterogeneous sources to generate faithful rationales. Unlike previous studies, our approach constructs planning data for complex questions from KGs, recognizing that patterns within KGs inherently represent multi-step plans. This data is utilized to enhance the planning capabilities of LLMs.
@@
Recent research on complex logic queries in KGs primarily focuses on first-order logical (FOL) queries that incorporate operations like conjunctions, disjunctions, negation, and existential quantifiers within incomplete KGs Hamilton et al. (2018); Ren et al. (2020); Ren and Leskovec (2020); Arakelyan et al. (2021); Chen et al. (2022); Xu et al. (2022). These works define diverse patterns to assess the capability of logical operations in vector spaces, specifically targeting logical forms rather than natural language. Nonetheless, their methodologies for pattern definition and extraction inspire our approach to deriving complex questions from KGs.
@@ 3.2 Construction of Planning Data
where $\{Q_{s_{n}}\}_{n=1}^{k}\text{ and }Q_{c}$ represent the resulting sub-questions and complex question respectively, $concat$ is string level concatenation. We use GPT-4 as $llm$ here. It is important to note that here the $llm$’s role is merely to transform the data format; the sub-questions and complex question still originate from the structure of the KG itself, without introducing any knowledge from the $llm$ in the task of question planning. The prompt we use can be found in Appendix B.1.
@@
where .$fill$ is a filling function of templates $\mathcal{T}_{pat}$. Inspired by Aksitov et al. (2023), we use a code-formatted input $x\text{ and output }y$ here (shown in “Input” and “Output” in Figure 2) to facilitate formatting and subsequent parsing and execution of the output plan (more details in Appendix B.2). In the end, we obtain 9000 training data entries $\mathcal{D}_{train}=\{x_{n},y_{n}\}_{n=1}^{9000}$, with 1000 entries for each pattern. We randomly select 100 items from the training sets for manual verification, with an accuracy rate of over 95%.
@@ 3.3 Fine-tuning and Inference of Planning LLMs
where $Ins$ and $De_{2}$ are the same as the contents in the Equation (3). It should be noted that in the multi-hop questions, the specific sub-questions in the second and third hops need to be constructed based on the answers to the previous hop’s sub-questions. Since our $P$ outputs all processes at once, the $\mathcal{M}_{p}$ cannot know the answers to the previous hop’s sub-questions when outputting the plans. Therefore, we will use a placeholder to replace the answer to the previous hop sub-questions, allowing the planning to proceed smoothly (as shown in Table 8, 9, 12, 13 in Appendix B.1). These placeholders will then be filled in during the subsequent parsing and execution process.
@@ 3.4 Plan Parsing and Execution
$\bullet$ When a step includes a “Get Answer” function, we’ll invoke an external QA LLM $\mathcal{M}_{QA}$ to get answers for a sub-question based on the retrieved information. The possible placeholders in sub-questions will be filled with previous answers. We ask QA LLM to organize answers in the form of a list (prompt is shown in Table 6 in Appendix B.3).
@@ 5 Experiment
Planning RAG HotPotQA 2WikiQA Bamboogle MuSiQue
@@
ICLPKG(GPT-3.5) 0.352 0.344 0.272 0.254
LPKG(CodeQwen) 0.338 0.356 0.256 0.266
LPKG(Llama3) 0.376 0.372 0.272 0.296
@@
HotPotQA 2WikiQA Bamboogle MuSiQue
@@ 5.1 Experimental Settings
We first conduct experiments on four conventional complex QA datasets: HotPotQA Yang et al. (2018), 2WikiMultiHopQA(2WikiQA) Ho et al. (2020), MuSiQue Trivedi et al. (2022), and Bamboogle Press et al. (2023). Among them, HotPotQA, 2WikiQA, and MuSiQue contain completed train sets, development sets, and test sets, while Bamboogle is a small dataset that only contains 125 test data. Similar to the previous method Shao et al. (2023); Aksitov et al. (2023), we respectively extract the first 500 entries from the development set of HotPotQA, 2WikiQA. For MuSiQue, we follow Press et al. (2023) to use only 2-hop questions in the development set. And for Bamboogle, we use all of its data as test data. Finally, we conduct testing on our benchmark CLQA-Wiki.
@@
Exact Match (EM) is set as an evaluation metric in HotPotQA, 2WikiQA, Bamboogle, and MuSiQue. While in CLQA-Wiki, we use Recall and Precision.
@@ 5.2 Results on Conventional Complex QA
At the same time, we also attempt to replace the fine-tuned $\mathcal{M}_{p}$ with GPT-3.5 while keeping other parts unchanged, denoted as “ICLPKG(GPT-3.5)” in Table 2. Results show that even though fine-tuned $\mathcal{M}_{p}$ CodeQwen (7B) and Llama3 (8B) have significantly fewer parameters than GPT-3.5 (more than 175B), they can maintain or even surpass GPT-3.5 in terms of planning ability. Next, we replace the $\mathcal{M}_{p}$ fine-tuned on KG-sourced data with their raw models, and the experimental results are shown in Table 3. It can be observed that after fine-tuning with planning data derived from the KG, both CodeQwen1.5-7B-Chat and Llama3-8B-Instruct show significant improvements in planning ability. In particular, CodeQwen1.5-7B-Chat, which is significantly inferior to GPT-3.5 across all datasets in planning ability before fine-tuning, exhibits a notable enhancement after fine-tuning on KG-based planning data, especially achieving better results than GPT-3.5 on 2WikiQA and MuSiQue. All these experimental phenomena fully demonstrate the efficacy of using KG-sourced planning data in improving the planning ability of the LLMs.
@@
To further validate the effectiveness of using planning data constructed from KG, we compare it with the normal distillation method. Specifically, we extracted 3000 questions each from the training sets of HotPotQA, 2WikiQA, and MuSiQue (9000 questions in total). Using the same input prompt with Equation (6), we obtain the planning process of these questions by invoking GPT-3.5. These planning data are then used to fine-tune CodeQwen1.5-7B-Chat which has relatively weaker planning capabilities, resulting in DLPKG(CodeQwen).
@@ Appendix A Case Study
Appendix B Prompt Content
B.1 Prompt of Verbalization
Table 7,8,9,10,11,12,13,14,15 shows the different prompts of verbalizing pattern instances to their natural language questions. The specific instance that needs to be verbalized will be added to the end of the prompt.
B.2 Prompt of Planning LLM

### v2

@@ Abstract
Junjie Wang1,2,511footnotemark: 1, Mingyang Chen3thanks:   Equal contribution., Binbin Hu2,5, Dan Yang2,5, Ziqi Liu2,5, Yue Shen2,5, Peng Wei2,5, Zhiqiang Zhang2,5, Jinjie Gu2,5, Jun Zhou2,5, Jeff Z. Pan4, Wen Zhang1,522footnotemark: 2, Huajun Chen1,5thanks:   Corresponding authors. 1Zhejiang University, 2Ant Group, 3Baichuan Inc., 4The University of Edinburgh 5Zhejiang University - Ant Group Joint Laboratory of Knowledge Graph {wangjj2018,zhang.wen,huajunsir}@zju.edu.cn, chenmingyang@baichuan-inc.com http://knowledge-representation.org/j.z.pan/ https://github.com/zjukg/LPKG
@@ 1 Introduction
The past few years have witnessed significant innovations in LLMs Ouyang et al. (2022); Touvron et al. (2023); Chowdhery et al. (2023); AI@Meta (2024). While LLMs excel in many natural language processing tasks, they still face challenges, particularly the smaller models, in handling complex question-answering (QA) tasks Press et al. (2023); Shao et al. (2023); Yao et al. (2022); Xiong et al. (2024a); Huang et al. (2024).
@@
To improve the performance of LLMs on complex QA tasks, past research has tried various methods: (1) Employing carefully designed prompt strategies to guide the model in reasoning, such as Chain of Thought (CoT) Kojima et al. (2022); Wei et al. (2022) and Tree of Thought (ToT) Yao et al. (2024) methods; (2) Utilizing retrieval techniques to obtain supplemental information from external knowledge source Lewis et al. (2020); Guu et al. (2020); (3) Combining prompt strategies with retrieval enhancements, as exemplified by methods like ReAct Yao et al. (2022) and Self-Ask Press et al. (2023). The third approach has garnered widespread research interest due to its integration of the advantages of the first two methods. The fundamental idea of this class of methods is to guide LLMs in breaking down a complex question into multiple simpler sub-questions and then use a retrieval-augmented generation (RAG) Huang et al. (2023, 2024) method to answer each sub-question, thereby deducing the answer to the original complex question. However, planning for complex questions is non-trivial, especially for smaller LLMs (with fewer than 10 billion parameters), which often require supervised fine-tuning Aksitov et al. (2023); Chen et al. (2023a); Qin et al. (2023).
@@
Knowledge Graphs (KGs) Pan et al. (2017b, a) usually store accurate knowledge in a structured way. We find that a KG pattern can be viewed as the abstract of a complex question, as shown in Figure 1, which reveals the connection between question planning and patterns. This opens up the possibility of constructing training data to enhance the planning capabilities of LLMs using KGs. Specifically, we start by grounding predefined patterns in an open-domain KG to extract numerous instances, which we then verbalize into complex questions and corresponding sub-questions in natural language. In this way, we effectively create a large number of accurate planning data for fine-tuning. Being fine-tuned with these planning data, LLMs’ capability of generating plans for complex questions is enhanced, resulting in better final answers by parsing and executing these plans. We refer to this innovative framework as Learning to Plan from Knowledge Graphs (LPKG).
@@ 2 Related Works
In the existing realm of LLMs, KGs are primarily utilized as sources of structured factual knowledge Pan et al. (2023). For example, Think-on-Graph Sun et al. (2023) extracts relevant triples from KGs to assist in QA. Reasoning on Graph (RoG) Luo et al. (2023) generates relation-based plans and retrieves corresponding paths from these graphs. While aiding in KGQA tasks where answers are directly sourced from KGs, these graphs also support rationale generation. Chain-of-Knowledge (CoK) Li et al. (2024) further leverages KGs along with other heterogeneous sources to generate faithful rationales. Unlike previous studies, our approach constructs planning data for complex questions from KGs, recognizing that patterns within KGs inherently represent multi-step plans. This data is utilized to enhance the planning capabilities of LLMs.
@@
Recent research on complex logic queries in KGs primarily focuses on first-order logical (FOL) queries that incorporate operations like conjunctions, disjunctions, negation, and existential quantifiers within incomplete KGs Hamilton et al. (2018); Ren et al. (2020); Ren and Leskovec (2020); Arakelyan et al. (2021); Chen et al. (2022); Xu et al. (2022); Xiong et al. (2024b); Wu et al. (2024). These works define diverse patterns to assess the capability of logical operations in vector spaces, specifically targeting logical forms rather than natural language. Nonetheless, their methodologies for pattern definition and extraction inspire our approach to deriving complex questions from KGs.
@@ 3.2 Construction of Planning Data
where $\{Q_{s_{n}}\}_{n=1}^{k}\text{ and }Q_{c}$ represent the resulting sub-questions and complex question respectively, $concat$ is string level concatenation. We use GPT-4 as $llm$ here. It is important to note that here the $llm$’s role is merely to transform the data format; the sub-questions and complex question still originate from the structure of the KG itself, without introducing any knowledge from the $llm$ in the task of question planning. The prompt we use can be found in Appendix C.1.
@@
where .$fill$ is a filling function of templates $\mathcal{T}_{pat}$. Inspired by Aksitov et al. (2023), we use a code-formatted input $x\text{ and output }y$ here (shown in “Input” and “Output” in Figure 2) to facilitate formatting and subsequent parsing and execution of the output plan (more details in Appendix C.2). In the end, we obtain 9000 training data entries $\mathcal{D}_{train}=\{x_{n},y_{n}\}_{n=1}^{9000}$, with 1000 entries for each pattern. We randomly select 100 items from the training sets for manual verification, with an accuracy rate of over 95%.
@@ 3.3 Fine-tuning and Inference of Planning LLMs
where $Ins$ and $De_{2}$ are the same as the contents in the Equation (3). It should be noted that in the multi-hop questions, the specific sub-questions in the second and third hops need to be constructed based on the answers to the previous hop’s sub-questions. Since our $P$ outputs all processes at once, the $\mathcal{M}_{p}$ cannot know the answers to the previous hop’s sub-questions when outputting the plans. Therefore, we will use a placeholder to replace the answer to the previous hop sub-questions, allowing the planning to proceed smoothly (as shown in Table 9, 10, 13, 14 in Appendix C.1). These placeholders will then be filled in during the subsequent parsing and execution process.
@@ 3.4 Plan Parsing and Execution
$\bullet$ When a step includes a “Get Answer” function, we’ll invoke an external QA LLM $\mathcal{M}_{QA}$ to get answers for a sub-question based on the retrieved information. The possible placeholders in sub-questions will be filled with previous answers. We ask QA LLM to organize answers in the form of a list (prompt is shown in Table 7 in Appendix C.3).
@@ 5 Experiment
Planning RAG HotPotQA 2WikiMQA Bamboogle MuSiQue
@@
ICLPKG(GPT-3.5) 0.352 0.344 0.296 0.254
LPKG(CodeQwen) 0.338 0.356 0.280 0.266
LPKG(Llama3) 0.376 0.372 0.304 0.296
@@
HotPotQA 2WikiMQA Bamboogle MuSiQue
@@ 5.1 Experimental Settings
We first conduct experiments on four conventional complex QA datasets: HotPotQA Yang et al. (2018), 2WikiMultiHopQA(2WikiMQA) Ho et al. (2020), MuSiQue Trivedi et al. (2022), and Bamboogle Press et al. (2023). Among them, HotPotQA, 2WikiMQA, and MuSiQue contain completed train sets, development sets, and test sets, while Bamboogle is a small dataset that only contains 125 test data. Similar to the previous method Shao et al. (2023); Aksitov et al. (2023), we respectively extract the first 500 entries from the development set of HotPotQA, 2WikiMQA. For MuSiQue, we follow Press et al. (2023) to use only 2-hop questions in the development set. And for Bamboogle, we use all of its data as test data. Finally, we conduct testing on our benchmark CLQA-Wiki.
@@
Exact Match (EM) is set as an evaluation metric in HotPotQA, 2WikiMQA, Bamboogle, and MuSiQue. While in CLQA-Wiki, we use Recall and Precision.
@@ 5.2 Results on Conventional Complex QA
At the same time, we also attempt to replace the fine-tuned $\mathcal{M}_{p}$ with GPT-3.5 while keeping other parts unchanged, denoted as “ICLPKG(GPT-3.5)” in Table 2. Results show that even though fine-tuned $\mathcal{M}_{p}$ CodeQwen (7B) and Llama3 (8B) have significantly fewer parameters than GPT-3.5 (more than 175B), they can maintain or even surpass GPT-3.5 in terms of planning ability. Next, we replace the $\mathcal{M}_{p}$ fine-tuned on KG-sourced data with their raw models, and the experimental results are shown in Table 3. It can be observed that after fine-tuning with planning data derived from the KG, both CodeQwen1.5-7B-Chat and Llama3-8B-Instruct show significant improvements in planning ability. In particular, CodeQwen1.5-7B-Chat, which is significantly inferior to GPT-3.5 across all datasets in planning ability before fine-tuning, exhibits a notable enhancement after fine-tuning on KG-based planning data, especially achieving better results than GPT-3.5 on 2WikiMQA and MuSiQue. All these experimental phenomena fully demonstrate the efficacy of using KG-sourced planning data in improving the planning ability of the LLMs.
@@
To further validate the effectiveness of using planning data constructed from KG, we compare it with the normal distillation method. Specifically, we extracted 3000 questions each from the training sets of HotPotQA, 2WikiMQA, and MuSiQue (9000 questions in total). Using the same input prompt with Equation (6), we obtain the planning process of these questions by invoking GPT-3.5. These planning data are then used to fine-tune CodeQwen1.5-7B-Chat which has relatively weaker planning capabilities, resulting in DLPKG(CodeQwen).
@@
Error Analysis
Planning Error Retrieval Error QA LLM Error
13 17 10
Table 5: Error analysis of LPKG(Llama3).
To gain a deeper understanding of the model’s performance, we conduct an error analysis of LPKG. Specifically, we extract 40 incorrect samples (10 per dataset) of LPKG(Llama3) and manually categorize the error cases into three types: planning error, retrieval error, and QA LLM error. As shown in Table 5, the performance of the retrieval model has the greatest impact. Among the 13 samples with planning errors, 10 of them are due to incorrect judgment of the type of questions, and 3 are due to incorrect expression of sub-questions. Future exploration directions can be based on improving the performance of the retriever model and enhancing the planning LLM’s ability to identify question types.
@@ 6 Conclusion
Acknowledgement
This work is founded by National Natural Science Foundation of China (NSFC62306276/NSFCU23B2055/NSFCU19B2027), Zhejiang Provincial Natural Science Foundation of China (No. LQ23F020017), Yongjiang Talent Introduction Programme (2022A-238-G), Chang Jiang Scholars Program (J2019032), Fundamental Research Funds for the Central Universities (226-2023-00138).
@@ Appendix A Case Study
Appendix B Prevention of Data Leakage
In the experiment, we used data from knowledge graph sources as training data, which may raise concerns about data leakage, specifically the overlap between the training data and the four multi-hop test sets(HotPotQA, 2WikiMQA, Bamboogle, MusiQue). We calculate the semantic similarity between the training and testing questions. We use the BGE-M3 model to embed the training and testing questions into high-dimensional vectors and calculate their cosine similarity. We found that the similarity between the testing and training questions does not exceed 0.8, and is generally below 0.6. This indicates that there is almost no high degree of overlap between the training data and the testing data.
In addition, when conducting experiments on our own test set CLQA-Wiki, we excluded questions similar to those in CLQA-Wiki from the training data (with similarity scores above 0.9), ensuring no high overlap between the training and testing data.
Appendix C Prompt Content
C.1 Prompt of Verbalization
Table 8,9,10,11,12,13,14,15,16 shows the different prompts of verbalizing pattern instances to their natural language questions. The specific instance that needs to be verbalized will be added to the end of the prompt.
C.2 Prompt of Planning LLM